Demarcus has $30, and Liu has $14. Demarcus is saving $4 per day, and Liu is saving $6 per day. After how many days will Demarcus and Liu have the same amount of money?
Answer by richwmiller(17219)

s=30+4d
s=14+6d
30+4d=14+6d
16=2d
8=d days